A particle is moving along the x-axis. Its acceleration is given by π‘Žπ‘₯ = 𝑐𝑑 βˆ’ 𝑑𝑑2

with c=2.0m/s3and d=0.24m/s4.The particle is at rest at the origin at t=0. Find position andΒ 

velocity as functions of time. What is the maximum velocity the particle reaches?



1
Expert's answer
2021-12-10T11:30:46-0500
"v=\int(ct-dt^2)\text dt=t^2-0.08t^3.\\
v_\text{ max}=23.1\text{ m/s}.\\
x=\int(t^2-0.08t^3)\text dt=\frac{t^3}3-0.02t^4."